OpenStudy (anonymous):

Provide an example that uses the elimination method to solve two linear equations.

OpenStudy (dumbcow):

4x - 2y = 10 3x + 4y = 13 get the coefficients in y column equal by multiplying top equation by 2 8x-4y=20 3x+4y=13 add each column 11x +0y = 33 solve for x 11x=33 x=33/11 =3 Then substitute 3 in for x in one of original equations 4x-2y =10 4(3) -2y =10 12-2y=10 -2y = -2 y=-2/-2 = 1 solution is x=3,y=1

OpenStudy (anonymous):

Thank you